OpenStudy (anonymous):

-(1)/(7)+(1)/(14)=(1)/(14)

OpenStudy (anonymous):

\[\frac{-1}{7} + \frac{1}{14} \neq \frac{1}{14}\]\[\frac{-2 + 1 }{14} = \frac{-1}{14}\] what's your question? ^_^

OpenStudy (anonymous):

adding thiis and they want you to simplfiy

OpenStudy (anonymous):

adding this and they want me to simpliy this

OpenStudy (anonymous):

what do you mean? don't you want to add the left side to check if it's equal to the right side?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but it's not equal you'll get the following:\[\frac{-1}{14} = \frac{1}{14}\] which are not equal

OpenStudy (anonymous):

am on a nother one no one no that one